OpenStudy (vishweshshrimali5):

solve 21÷ 3 (10 - 3) - 20 + 1 = ?

OpenStudy (vishweshshrimali5):

solve 21\(\div\) 3 (10 - 3) - 20 + 1 = ?

OpenStudy (vishweshshrimali5):

@Kainui

OpenStudy (lgbasallote):

i smell a trickedy-trick

OpenStudy (vishweshshrimali5):

Not a trick @lgbasallote just trying to clear all MISCONCEPTIONS

OpenStudy (vishweshshrimali5):

comeon try it

OpenStudy (lgbasallote):

i like working from left to right so i will \[21 \div 3(10 - 3) - 20 + 1\] \[\implies 7(10-3) - 20 + 1\] \[\implies 7(7) - 20 +1\] \[\implies 49 - 20 + 1\] \[\implies 29 + 1\] \[\implies 30\]

OpenStudy (vishweshshrimali5):

Gud and thats what I did in the previous question

OpenStudy (vishweshshrimali5):

And now tell me why didn't u take it as \[\large \cfrac{21}{3(10-3)}\]

OpenStudy (vishweshshrimali5):

@lgbasallote

OpenStudy (lgbasallote):

\[\frac{21}{3(10-3)} \implies 21 \div \left[3(10-3)\right]\] that would be the proper way to write it

OpenStudy (vishweshshrimali5):

so in other words u used VBODMAS rule ......... right ?

OpenStudy (lgbasallote):

it's actually the left to right rule...

OpenStudy (lgbasallote):

you should always operate from left to right

OpenStudy (vishweshshrimali5):

really ??????? now stick to ur rule.. ok

OpenStudy (vishweshshrimali5):

what is \[\large \cfrac{(4+4+4)\div 4}{6+6+6\div 6}\]

OpenStudy (lgbasallote):

but you should follow PEMDAS always

OpenStudy (lgbasallote):

or whatever acronym you want

OpenStudy (lgbasallote):

\[\implies \frac{ 12 \div 4}{6+6 + 1} \implies \frac{3}{13}\]

OpenStudy (vishweshshrimali5):

done ....

OpenStudy (vishweshshrimali5):

where has ur left to right rule gone now ?

OpenStudy (lgbasallote):

i operated the numerator from left to right

OpenStudy (lgbasallote):

then operate the denominator left to right

OpenStudy (lgbasallote):

then divide

OpenStudy (lgbasallote):

satisfied yet @vishweshshrimali5 ?

OpenStudy (vishweshshrimali5):

really @lgbasallote if u would have went from left to right then first u would have added 6 + 6 + 6 and then would have divided by 6 i.e. u would have got 3 in denominator

OpenStudy (vishweshshrimali5):

what u did is not a left to right rule but the VBODMAS rule

OpenStudy (lgbasallote):

lol didnt you read what i wrote...you still follow PEMDAS or whatever acronym you want

OpenStudy (lgbasallote):

tsk tsk. read read.

OpenStudy (vishweshshrimali5):

now what is tsk

OpenStudy (lgbasallote):

PEMDAS or whatever acronym becomes unreliable in the problem you wrore unless you go from left to right.

OpenStudy (lgbasallote):

that's the additional rule (probably not taught to a lot of students)

OpenStudy (lgbasallote):

tsk tsk is a sound of disappointment

OpenStudy (vishweshshrimali5):

ohh

OpenStudy (vishweshshrimali5):

well leave everything I just want to say is that we should stay fixed to predecided rules

OpenStudy (lgbasallote):

i did stick to the :"predecided rules" you dont listen well do you =))

OpenStudy (vishweshshrimali5):

Check out this....... ohh so left to right is a predecided rule ......... don't say yes !!!

OpenStudy (lgbasallote):

the left to right rule is a rule in PEMDAS or whatever acronym

OpenStudy (lgbasallote):

see how that pdf solves from left to right

OpenStudy (lgbasallote):

\[21 \div 3(10-3) - 20 + 1\] if you dont solve from left to right you end up with \[21 \div 21 - 20 + 1\]

OpenStudy (lgbasallote):

got the point of the left to right rule im talking about?

OpenStudy (lgbasallote):

im not saying it is a rule like PEMDAS..the left to right rule is an ADDITIONAL rule

OpenStudy (kainui):

\[\frac{ 21 }{ 3 }(10 - 3) - 20 + 1\] or\[\frac{ 21 }{ 3(10 - 3) } - 20 + 1\]What is the right equation? I believe the bottom one is correct because the 3 is distributed to the numbers, not (1/3). However, if the original equation was: 21÷ 3 * (10 - 3) -20 + 1 Then it would be the top one since the 3 is no longer a part of the parentheses. When a number is next to another number like that, it kind of shows it was factored out. Let's replace the (10-3) with an x and disregard the rest. 21÷3x Now does that become \[\frac{ 21 }{ 7x }\] or\[3x\] Hint: it's definitely not 3x. I've done enough algebra in my life to know this.

OpenStudy (vishweshshrimali5):

I am sorry to tell you know that you still not have done enough algebra to know this. Your answer is WRONG.

OpenStudy (vishweshshrimali5):

@Kainui

mathslover (mathslover):

ok wait

OpenStudy (vishweshshrimali5):

@lgbasallote correct

mathslover (mathslover):

solve 21÷ 3 (10 - 3) - 20 + 1 = ? \[\large{21 \div 3*7-20+1=1-20+1=-18}\]

OpenStudy (vishweshshrimali5):

And same is with u @mathslover

mathslover (mathslover):

VBODMAS here BODMAS is applied.. B ---> brackets O ---> of D ----> div M ----> multiplication A-----> addition S-----> subtraction

OpenStudy (vishweshshrimali5):

I can't believe that you don't know anything about VBODMAS rule

OpenStudy (vishweshshrimali5):

Really @mathslover

OpenStudy (vishweshshrimali5):

I think you have forgotten how to apply the rules @mathslover

OpenStudy (vishweshshrimali5):

u were supposed to divide first ........ i.e. \[21\large {\div}3\]

mathslover (mathslover):

no ... brackets first @vishweshshrimali5

mathslover (mathslover):

solving brackets first then dividing is right

OpenStudy (vishweshshrimali5):

I am talking about \(21 \div 3\) part only

OpenStudy (vishweshshrimali5):

ok so

mathslover (mathslover):

you mean to say that: \[\large{\frac{21}{3} *(10-3) -20+1}\] ?

OpenStudy (kainui):

Lol I mean this whole thing is completely ridiculous. In actual practice no one ever writes things as ambiguous as that. It just doesn't happen. It's either in one of the two forms I wrote, I never see the ÷ sign when someone means that it's a fraction.

OpenStudy (vishweshshrimali5):